Two charges Q1=5nc and Q2=-7nc are separated by 8 cm. Calculate the electric potential at point A


1
Expert's answer
2022-04-21T17:44:45-0400

The electric potential at any point "A" is given as follows:


"\varphi_A = k\dfrac{Q_1}{r_1} + k\dfrac{Q_2}{r_2}"

where "Q_1=5nC, Q_2 = -7nC" are the charges, "r_1, r_2" are the distances to point "A" from the first and the second charge respectively, and "k\approx 9\times 10^9N\cdot m^2/C^2" is the Coulomb's constant. The particular value of "\varphi _A" depends on the location of point "A".